# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v3

@@ A.1 Lemmas
$\sup_{i,t}\mathbb{E}\left(|\partial_{\phi_{1}\ldots\phi_{S}}l_{it}\left(\theta,\phi\right)|\mathds{1}\{|\partial_{\phi_{1}\ldots\phi_{S}}l_{it}\left(\theta,\phi\right)|>a\}\right)\rightarrow 0$
@@
$\sqrt{NT}\mathbb{E}\left\{\big{[}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)+s_{\breve{\gamma}}^{\prime}\left(\theta\right)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)s_{\breve{\gamma}}(\theta)\big{]}\cdot o_{\mathbb{P}}(1)\right\}\rightarrow 0,$
as $N,T\rightarrow\infty.$ This completes the proof of Lemma 1.
@@
Under Assumption 1, $r^{s}(\theta,a)$ defined by (2.3) satisfies
@@
as $N$ and $T\rightarrow\infty$, where $a=\left(a_{1},\ldots,a_{\left(N+T-2\right)K}\right)^{\prime}$ for each component $a_{j}\in(0,1)$.
Proof. In the proof, we retain the notation and setup from Lemma 1. From equations (2.3) and (2.5), it is evident that the dominant term in the expansion of $\widehat{\psi}(\theta)-\psi\left(\theta\right)$ is $[-{H}(\theta)]^{-1}s(\theta).$ Specifically,
@@
$=-\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\end{pmatrix}+[-\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\cdot o_{\mathbb{P}}(1)$ (A.6)
@@
where $o_{\mathbb{P}}(1)$ is uniform in both $i$ and $t.$ Let $u_{i}^{\breve{\alpha}}(\theta)$ denote the $i$-th element of $H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)$ for $i=1,\ldots,N-1.$ It follows that
$N\sqrt{T}u_{i}^{\breve{\alpha}}(\theta)=N\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\big{[}\partial_{\alpha_{j}}\widetilde{l}_{j,t}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\big{]}\bigg{)}.$ (A.7)
@@
Using the result from Miller (1981), we know that $\sup_{i\geq 1}h_{i,i}^{\ast\breve{\alpha}}(\theta)=O(1)$ and $\sup_{i\neq j}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in $\theta,$ and it also implies that $\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta.$ Since $\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta))=O_{\mathbb{P}}(1)$ uniformly over $i$ by Assumptions 1 and 1, we obtain
@@
$N\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\partial_{\alpha_{j}}\widetilde{l}_{j,t}(\theta,\phi(\theta))\bigg{)}=O_{\mathbb{P}}(1)$
and
$N\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\cdot\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\bigg{)}=O_{\mathbb{P}}(1).$
@@
Therefore, $N\sqrt{T}u_{i}^{\breve{\alpha}}(\theta)=O_{\mathbb{P}}(1)$ uniformly for $i.$ Let $u_{i}^{\breve{\alpha}\breve{\gamma}}(\theta)$ denote the $i$-th element of $H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\breve{\gamma}}(\theta)$ for $i=1,\ldots,N-1.$ Given that $\sup_{\theta\in\Theta}\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\|_{\max}=O(1),$ we have
@@
$\sqrt{N}Tu_{i}^{\breve{\alpha}\breve{\gamma}}(\theta)=\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{(}\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\big{[}\partial_{\gamma_{t}}\widetilde{l}_{j,t}(\theta,\phi(\theta))-\partial_{\gamma_{T}}\widetilde{l}_{j,T}(\theta,\phi(\theta))\big{]}\bigg{)}.$ (A.8)
@@
Since $\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)=O(1),$ which follows from $\mathbb{E}[\partial_{\breve{\alpha}\breve{\gamma}^{\prime}}l^{*}(\theta,\psi)|_{\psi=\psi(\theta)}]=\partial_{\breve{\gamma}^{\prime}}\mathbb{E}[\partial_{\breve{\alpha}}l^{*}(\theta,\psi)|_{\psi=\psi(\theta)}]=0,$ it holds that
@@
$\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{(}\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\partial_{\gamma_{t}}\widetilde{l}_{j,t}(\theta,\phi(\theta))\bigg{)}=O_{\mathbb{P}}(1),$
and
$\bigg{(}\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\bigg{(}\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\partial_{\gamma_{T}}\widetilde{l}_{j,T}(\theta,\phi(\theta))\bigg{)}=O_{\mathbb{P}}(1),$
@@
$\sup_{\theta\in\Theta}\|H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\|_{\max}=O_{\mathbb{P}}\big{(}1/(N\sqrt{T})\big{)},\quad\sup_{\theta\in\Theta}\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\|_{\max}=O_{\mathbb{P}}\big{(}1/(\sqrt{N}T)\big{)}.$ (A.9)
Similarly, we can show that
$\sup_{\theta\in\Theta}\|H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\|_{\max}=O_{\mathbb{P}}\big{(}1/(\sqrt{N}T)\big{)},\quad\sup_{\theta\in\Theta}\|H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\|_{\max}=O_{\mathbb{P}}\big{(}1/(N\sqrt{T})\big{)}.$ (A.10)
@@
$\left\|[-{H}(\theta)]^{-1}s(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\big{(}1/(NT)^{3/4}\big{)},$
@@
$\left\|\widehat{\psi}(\theta)-\psi(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\big{(}1/(NT)^{3/4}\big{)}.$ (A.11)
@@
$\partial_{\alpha_{i}}H(\theta)=\begin{pmatrix}\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}\left(\theta\right),&\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)\\ \partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\alpha}^{\prime}}\left(\theta\right),&\partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)\end{pmatrix}.$
In particular, let $\text{diag}\{a_{1},\ldots,a_{n}\}$ denote a diagonal matrix with elements $a_{1},\ldots,a_{n}$ on its diagonal. We then have
@@
$N\cdot\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}\left(\theta\right)=$ $\frac{1}{T}\sum_{t=1}^{T}\text{diag}\big{\{}0,\ldots,0,\partial_{\alpha_{i}\alpha_{i}\alpha_{i}}l_{i,t}(\theta,\phi)|_{\phi=\phi(\theta)},0,\ldots,0\big{\}}$
@@
$\quad-\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}\,\iota_{N-1}\iota_{N-1}^{\prime},$
@@
$NT\cdot\partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}\left(\theta\right)=$ $\text{diag}\left\{\partial_{\gamma_{1}\gamma_{1}\alpha_{i}}l_{i,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\gamma_{T-1}\gamma_{T-1}\alpha_{i}}l_{i,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\right\}$
$\quad+\partial_{\gamma_{T}\gamma_{T}\alpha_{i}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\,\iota_{N-1}\iota_{N-1}^{\prime},$
@@
$NT\cdot\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}\left(\theta\right)=$ $\big{(}\partial_{\alpha_{i}\alpha_{i}\gamma_{1}}l_{i,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\alpha_{i}\alpha_{i}\gamma_{T-1}}l_{i,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\big{)}\otimes e_{i}$
@@
$\quad+\big{(}\partial_{\alpha_{N}\alpha_{N}\gamma_{1}}l_{N,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\alpha_{N}\alpha_{N}\gamma_{T-1}}l_{N,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\big{)}\otimes\iota_{N-1},$
@@
where $e_{i}$ is an $(N-1)\times 1$ vector with the $i$-th component equal to 1 and all other elements equal to 0. Using (A.11), for each $j=1,\ldots,N-1,$ we write
$\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
$=$ $\sup_{a_{j}\in(0,1)}\frac{1}{2}\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}G_{\psi_{j}}(\theta,a_{j})\big{(}\widehat{\psi}(\theta)-\psi\left(\theta\right)\big{)}$
@@
$=$ $\frac{1}{2}\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\end{pmatrix}^{\prime}\partial_{\alpha_{i}}H(\theta)\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\end{pmatrix}\big{(}1+o_{\mathbb{P}}(1)\big{)}$
@@
$=$ $\frac{1}{2}\left\{\big{(}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}^{\prime}\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)\big{(}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}\right.$
$\qquad+\big{(}H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}^{\prime}\partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)\big{(}H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}$
@@
$\qquad\left.+2\big{(}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}^{\prime}\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)\big{(}H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}\right\}\big{(}1+o_{\mathbb{P}}(1)\big{)}$
=:absent:=:= : $\frac{1}{2}\left\{C_{NT,1}(\theta)+C_{NT,2}(\theta)+C_{NT,3}(\theta)\right\}\big{(}1+o_{\mathbb{P}}(1)\big{)}.$ (A.12)

### v4

@@ A Proofs
$\sup_{i,t}\mathbb{E}\left(|\partial_{\phi_{1}\ldots\phi_{S}}l_{i,t}\left(\theta,\phi\right)|\mathds{1}\{|\partial_{\phi_{1}\ldots\phi_{S}}l_{i,t}\left(\theta,\phi\right)|>a\}\right)\rightarrow 0$
@@
$\sqrt{NT}\mathbb{E}\left\{\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+\widetilde{s}_{\gamma}^{\prime}\left(\theta\right)H_{\gamma\gamma^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\gamma}(\theta)\right]\cdot o_{\mathbb{P}}(1)\right\}\rightarrow 0,$
as $N,T\rightarrow\infty$. This completes the proof.
The next lemma shows that the remainder term in Equation (2.13), $r^{s}\left(\theta\right):=r^{s}\left(\theta,a\right)$ is $O_{\mathbb{P}}\left(1/\left(NT\right)\right)$ uniformly, where $r^{s}\left(\theta,a\right)$ is defined as [^10] Feng et al. (2014) point out that the expansion in (2.13) needs to be done component-by-component. This leads to a different $a_{j}$ for each component of $r^{s}\left(\theta,a\right)$.
$r^{s}\left(\theta,a\right)$ $:=\left(r_{\psi_{1}}^{s}\left(\theta,a_{1}\right),\ldots,r_{\psi_{(N+T-2)K}}^{s}(\theta,a_{(N+T-2)K})\right)^{\prime},$ (A.6)
$r_{\psi_{j}}^{s}\left(\theta,a_{j}\right)$ $:=\frac{1}{2}\left(\widehat{\psi}\left(\theta\right)-\psi(\theta)\right)^{\prime}G_{\psi_{j}}(\theta,a_{j})\left(\widehat{\psi}(\theta)-\psi(\theta)\right),$
with $a:=(a_{1},\ldots,a_{(N+T-2)K})^{\prime}$ for each $a_{j}\in(0,1)$, and $\psi_{j}$ denotes the $j$-th component of $\psi$. Here $G_{\psi_{j}}(\theta,a_{j})$ is the derivative of each element of $H(\theta,\psi)$ with respect to $\psi_{j}$, evaluated at the point $a_{j}\widehat{\psi}\left(\theta\right)+\left(1-a_{j}\right)\psi\left(\theta\right)$.
@@
Under Assumption 1,
@@
as $N$ and $T\rightarrow\infty$, where $\sup_{a}$ is taken over $a_{j}\in(0,1)$ for all $j=1,\ldots,(N+T-2)K$.
Proof. In the proof, we retain the notation and setup from Lemma A.1. From (2.13), it is evident that the dominant term in the expansion of $\widehat{\psi}(\theta)-\psi\left(\theta\right)$ is $[-{H}(\theta)]^{-1}s(\theta)$. Specifically,
@@
$=-\begin{pmatrix}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\\ H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\end{pmatrix}+[-\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\cdot o_{\mathbb{P}}(1)$ (A.7)
@@
where $o_{\mathbb{P}}(1)$ is uniform in both $i$ and $t$. Let $u_{i}^{\alpha}(\theta)$ denote the $i$-th element of $H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)$. It follows that
$N\sqrt{T}u_{i}^{\alpha}(\theta)=N\sum_{j=1}^{N-1}\left[h_{i,j}^{\ast\alpha}(\theta)\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\left[\widetilde{s}_{j,t}^{\alpha}(\theta)-\widetilde{s}_{N,t}^{\alpha}(\theta)\right]\right)\right].$ (A.8)
@@
Using the result from Miller (1981), we know that $\sup_{i\geq 1}h_{i,i}^{\ast\alpha}(\theta)=O(1)$ and $\sup_{i\neq j}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in $\theta$, implying $\sum_{j=1}^{N-1}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta$. Since $1/\sqrt{T}\sum_{t=1}^{T}\widetilde{s}_{i,t}^{\alpha}(\theta)=O_{\mathbb{P}}(1)$ uniformly over $i$ by Assumptions 1 and 1, we obtain
@@
$N\sum_{j=1}^{N-1}\left[h_{i,j}^{\ast\alpha}(\theta)\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\widetilde{s}_{j,t}^{\alpha}(\theta)\right)\right]=O_{\mathbb{P}}(1),\qquad N\left(\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\right)\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\widetilde{s}_{N,t}^{\alpha}(\theta)\right)=O_{\mathbb{P}}(1).$
@@
Therefore, $N\sqrt{T}u_{i}^{\alpha}(\theta)=O_{\mathbb{P}}(1)$ uniformly over $i$. Let $u_{i}^{\alpha\gamma}(\theta)$ denote the $i$-th element of $H_{\alpha\gamma^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\gamma}(\theta)$. Given that $\sup_{\theta\in\Theta}\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\|_{\max}=O(1)$, we have
@@
$T\sqrt{N}u_{i}^{\alpha\gamma}(\theta)=\sum_{t=1}^{T-1}\left[h_{i,t}^{\ast\alpha\gamma}(\theta)\left(\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\left[\widetilde{s}_{j,t}^{\gamma}(\theta)-\widetilde{s}_{j,T}^{\gamma}(\theta)\right]\right)\right].$ (A.9)
@@
Since $\sum_{t=1}^{T-1}h_{i,t}^{\ast\alpha\gamma}(\theta)=O(1)$, which follows from $\mathbb{E}[\partial_{\breve{\alpha}\breve{\gamma}^{\prime}}\ell(\theta,\psi)|_{\psi=\psi(\theta)}]=\partial_{\breve{\gamma}^{\prime}}\mathbb{E}[\partial_{\breve{\alpha}}\ell(\theta,\psi)|_{\psi=\psi(\theta)}]=0$, for $\breve{\alpha}:=\left(\alpha_{1}^{\prime},\ldots,\alpha_{N-1}^{\prime}\right)^{\prime}$ and $\breve{\gamma}:=\left(\gamma_{1}^{\prime},\ldots,\gamma_{T-1}^{\prime}\right)^{\prime}$, it holds that
@@
$\sum_{t=1}^{T-1}\left[h_{i,t}^{\ast\alpha\gamma}(\theta)\left(\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\widetilde{s}_{j,t}^{\gamma}(\theta)\right)\right]=O_{\mathbb{P}}(1),\qquad\left(\sum_{t=1}^{T-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\left(\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\widetilde{s}_{j,T}^{\gamma}(\theta)\right)=O_{\mathbb{P}}(1)$
@@
$\sup_{\theta\in\Theta}\|H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(N\sqrt{T})\right),\qquad\sup_{\theta\in\Theta}\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(T\sqrt{N})\right).$ (A.10)
Following a similar proof, we also have
$\sup_{\theta\in\Theta}\|H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(T\sqrt{N})\right),\qquad\sup_{\theta\in\Theta}\|H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(N\sqrt{T})\right).$ (A.11)
@@
$\left\|[-{H}(\theta)]^{-1}s(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\left(1/(NT)^{3/4}\right),$
@@
$\left\|\widehat{\psi}(\theta)-\psi(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\left(1/(NT)^{3/4}\right).$ (A.12)
@@
$\partial_{\alpha_{i}}H(\theta)=\begin{pmatrix}\partial_{\alpha_{i}}{H}_{\alpha\alpha^{\prime}}\left(\theta\right),&\partial_{\alpha_{i}}{H}_{\alpha\gamma^{\prime}}(\theta)\\ \partial_{\alpha_{i}}{H}_{\gamma\alpha^{\prime}}\left(\theta\right),&\partial_{\alpha_{i}}{H}_{\gamma\gamma^{\prime}}(\theta)\end{pmatrix}.$
We then have
@@
$N\cdot\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}\left(\theta\right)$ $=\frac{1}{T}\sum_{t=1}^{T}\operatorname*{diag}\left(0,\ldots,0,\partial_{\alpha_{i}\alpha_{i}\alpha_{i}}l_{i,t}(\theta,\phi)|_{\phi=\phi(\theta)},0,\ldots,0\right)$
@@
$\quad\quad-\left(\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\right)\iota_{N-1}\iota_{N-1}^{\prime},$
@@
$NT\cdot\partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}\left(\theta\right)$ $=\operatorname*{diag}\left\{\partial_{\gamma_{1}\gamma_{1}\alpha_{i}}l_{i,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\gamma_{T-1}\gamma_{T-1}\alpha_{i}}l_{i,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\right\}$
$\quad\quad+\partial_{\gamma_{T}\gamma_{T}\alpha_{i}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\,\iota_{N-1}\iota_{N-1}^{\prime},$
@@
$NT\cdot\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}\left(\theta\right)$ $=\left(\partial_{\alpha_{i}\alpha_{i}\gamma_{1}}l_{i,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\alpha_{i}\alpha_{i}\gamma_{T-1}}l_{i,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\right)\otimes e_{i}$
@@
$\quad\quad+\left(\partial_{\alpha_{N}\alpha_{N}\gamma_{1}}l_{N,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\alpha_{N}\alpha_{N}\gamma_{T-1}}l_{N,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\right)\otimes\iota_{N-1},$
@@
where $e_{i}$ is an $(N-1)\times 1$ vector with the $i$-th component equal to $1$ and all other elements equal to 00. Using (A.12), for each $j=1,\ldots,N-1$, we write
$\quad\quad\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
$=\sup_{a_{j}\in(0,1)}\frac{1}{2}\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}G_{\psi_{j}}(\theta,a_{j})\left(\widehat{\psi}(\theta)-\psi\left(\theta\right)\right)$
@@
$=\frac{1}{2}\begin{pmatrix}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\\ H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\end{pmatrix}^{\prime}\partial_{\alpha_{i}}H(\theta)\begin{pmatrix}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\\ H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\end{pmatrix}\left[1+o_{\mathbb{P}}(1)\right]$
@@
$=\frac{1}{2}\left\{\left(H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)^{\prime}\partial_{\alpha_{i}}{H}_{\alpha\alpha^{\prime}}(\theta)\left(H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)\right.$
@@
$\quad\quad+\left(H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)^{\prime}\partial_{\alpha_{i}}{H}_{\gamma\gamma^{\prime}}(\theta)\left(H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)$
$\quad\quad\left.+2\left(H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)^{\prime}\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)\left(H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)\right\}\left[1+o_{\mathbb{P}}(1)\right]$
$=:\frac{1}{2}\left\{C_{NT,1}(\theta)+C_{NT,2}(\theta)+C_{NT,3}(\theta)\right\}\left[1+o_{\mathbb{P}}(1)\right].$ (A.13)